

********************************************************************
UARS WEEKLY STATUS REPORT - 16 APR 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 70% of
each day's data amounting to nearly 780 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset, and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 April 2004 (Orbit 68728, GMT Day 094) through
09 April 2004 (Orbit 68883, GMT Day 100).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 094/1345 68737 POWER ON (PART 1)
PEM 095/1742 68754 AXIS 1 & 2 ON (START)
The beta angle ranged between the angles of 20.3 degrees to 47.9
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 68 degrees on 17 April 2004 (DOY 108).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.95 to 1.27 degrees C. Additional
battery performance data for each battery on 02 April and 09 April
2004 are provided in Appendix D.
The clock error ranged from 10.7 to -2.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HALOE = 02
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT transitioned to the 3-TDRS mode of operations on 17 March
2004. The FOT is now taking 35-38 events per day (up from 26-28 per
day under 2-TDRS operations) and collecting data for an additional 18-
20% more of each orbit. So far no problems have been found with this
mode and all subsystems remain nominal.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 07 May 2004 (DOY 128). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 April 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68736 094/1311 04-103 OBC/CLK NOOP XMTR ON CMD (PASS DELETED)
68737 094/1345 04-104 HALOE POWER ON (PART 1)
68740 094/1925 04-105 HALOE POWER ON (COMPLETE)
68752 095/1342 04-106 PEM AXIS 1 & 2 ON / OFF
68752 095/1424 04-106 PEM AXIS 1 & 2 ON (START)
68753 095/1453 04-106 PEM AXIS 2 ON - 1 OFF (COMPLETE)
68753 095/1603 04-107 PEM AXIS 2 OFF (UNUSUAL PWR VALUES)
68754 095/1711 04-106 PEM AXIS 1 & 2 ON (START)
68754 095/1742 04-106 PEM AXIS 1 & 2 ON (COMPLETE)
68827 100/1326 04-108 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 April 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 April 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68827 100/1338 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 09 April 2004
02 April 2004 - GMT Day 093 Beta = 18.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-22.4 +4.50/+3.23 +28.4/-12.8 26.1 1.03 19.3
3 +22.4/-33.6 +3.23/+2.28 +27.2/-12.4 26.1 1.01 19.1
09 April 2004 - GMT Day 100 Beta = 47.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-044.8 +4.82/+3.87 +25.6/-14.4 25.8 1.03 19.6
3 +11.2/-123.2 +3.87/+2.60 +24.8/-14.4 25.8 1.02 19.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 09 April 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-017 FOT 68744 095/0144 35517 Data drop out/11 minutes
04-018 FOT 68752 095/1342 N/A COP to INIT during commanding
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-014 FOT See List Mar 04 30832 Generic Late Acquisition
04-016 FOT 68708 092/1613 35490 Drop out due to HW @ WSGT
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-017 FOT 68744 095/0144 35517 Data drop out/11 minutes
04-018 FOT 68752 095/1342 N/A COP to INIT during commanding